# CrystalFormer-RL: Reinforcement Fine-Tuning for Materials Design

### v1

@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Many of the generated samples are present in the Alex-20 dataset which is in line with the S.U.N ratio 5.1% of the base model on space group No.225. This nevertheless demonstrates that the reinforcement fine-tuning procedure can guide the generative process toward discovering candidates within the pretrained dataset with no property label. Notably, 11 materials in Table LABEL:table:dielectric_samples are absent from the Alex-20 dataset, highlighting the ability of RL to guide the generative process toward novel samples beyond the pretrained dataset. We compare the prediction results of the generated samples with that of corresponding equilibrium structures from the Alex-20 dataset in Table LABEL:table:alex_dielectric_samples. We observe that the generated materials are close to the equilibrium structure in the Alex-20 dataset, with similar predicted FoM values.
@@ IV Discussions
Reinforcement fine-tuning of materials generative model is useful for an even broader class of materials properties such as the thermoelectric figure of merit, thermal conductivity, and superconducting critical temperature. In addition to a foundational generative model such as the CrystalFormer for the materials prior $p(x)$, the framework requires a reliable ML prediction model for the likelihood $p(y|x)$, which is a research topic with rapid progress both on generalizability and accuracy, see, e.g. [53, 54, 55].
@@
It is often easier to discriminate the utility of a given material than to generate a material with the desired utility. Reinforcement fine-tuning of a materials generative model using discriminative feedback offers a way to bridge such discrimination-generation gap. RL on the autoregressive transformer such as CrystalFormer [5] is particularly suitable given a variety of post-training techniques available for LLM. Although the RL technique is readily applicable to materials language models [56, 57, 4], we believe finetuned CrystalFormer is more advantageous because it cherishes the fundamental inductive bias of crystalline materials such as space group symmetry and periodicity. Furthermore, many of these RL methods can be extended to the material generative models based on diffusion. The diffusion model is typically fine-tuned through the incorporation of auxiliary networks to condition generation on external inputs [58], while RL eliminates the necessity for these additional network components. For example, Refs. [59, 60] employ the RL algorithm to fine-tune the pre-trained text-to-image model, leading to improved sample quality and diversity. Nevertheless, applying RL with KL regularization to the diffusion model is challenging due to the need for precise log-likelihood calculations [61].
@@
It is also interesting to compare active learning and RL in the context of generative materials discovery. Active learning [62, 63] selects crystal samples back to the generative training dataset, which is similar to the SFT approach discussed in Sec. III.1. The active learning procedure optimizes the same forward KL divergence between the empirical distribution of the training dataset and the generative model. In contrast, RL minimizes the reverse KL divergence [25] between the generative model and the target distribution Eq. (2). There are evidences that the RL approach delivers better performance compared to the active learning approach in the contexts of LLM fine-tuning [64]. Consequently, the primary advantage of RL over active learning is its ability to directly incorporate task-specific reward signals, enabling the model to generate outputs that more effectively meet practical application requirements. Moreover, we suspect the difference in the mode covering versus mode seeking in the active learning versus RL may also be a reason for the difference in the performance [65, 26].
@@
The stability verifications presented in Fig. 3 and Fig. 6(b) are based on the energy above the convex hull evaluated with an MLIP. This reveals a fundamental challenge in the application of RL in materials science: what is the verifiable ground truth reward signal. Given the intricacies involved in material synthesis and characterization [67], such a challenge is particularly concerning as faulty reward functions [68] which fails to align with the final goal will suggest structures with high reward but fail for the final experiment or even intermediate DFT verifications. We believe the design of reward functions that are both actionable and well-aligned with experimental outcomes, with the insights of human experts, is crucial for ensuring the reliability of the present and many related efforts.
@@
This study focuses on how generative models can be enhanced and specialized through the integration of property prediction models. However, the reverse question: how property prediction models can be improved by leveraging generative models—is also an intriguing question worth future investigation. For example, the Orb models employ a two-stage training framework that integrates generative and supervised learning to achieve state-of-the-art performance [13]. Initially, they are pre-trained as denoising diffusion models on datasets of stable materials. The pre-trained model then serves as the initialization for an MLIP, which is subsequently fine-tuned in a supervised manner to predict the energy, forces, and stress of materials. This approach exemplifies how generative pre-training can enhance the performance of discriminative models such as MLIP. Extending this idea further, Ref. [12] has demonstrated that fine-tuning pretrained MLIP also gives better performance for property prediction models.
@@ Acknowledgements.
We thank Han Wang, Linfeng Zhang, Jian Lv, Hongteng Xu, Shigang Ou, Xiaoshan Luo, Mingzhang Yang, Tianping Ying, and Ling Lu for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, and No. T2121001, National Key Projects for Research and Development of China Grant. No. 2021YFA1400400, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000.
@@ Appendix A The CrystalFormer model card
The base model was trained for 5500 epochs with a total batch size of 8000 over eight A100 GPUs using the Adam optimizer [73]. The learning rate was set to $48e-$ 4.
The reinforcement fine-tuning was performed for 100 - 200 steps, depending on the task. For the reinforcement fine-tuning, we use a total batch size of 1000 over two A100 GPUs and the Adam optimizer with a learning rate of $41e-$ 4. The pre-training phase required approximately 1000 GPU hours, while the reinforcement fine-tuning phase took around 12 to 24 GPU hours.
For the supervised fine-tuning (SFT) of the CrystalFormer, we generate 10,000 crystal samples and relax them using the Orb models. Only those stable samples with $E_{\mathrm{hull}}<0.1$ eV/atom were retained for SFT. The SFT was performed for 55 epochs when validation loss stopped improving, with a batch size of 100 using the Adam optimizer with a learning rate of $51e-$ 5 on a single A100 GPU.
@@
        Hidden layer dimension for the composite type of the first atom         256
@@
        RL regularization coefficient $\tau$         0.1         15 for FoM reward
@@ Appendix B Training dataset
Figure S3: Composition and structure validity of samples generated by the pretrained CrystalFormer using the MP-20 and Alex-20 dataset, as well as RL fine-tuned model.
@@ Appendix C Details of reinforcement fine-tuning implementation
where $\gamma(\theta)=\frac{p_{\theta}(x)}{p_{\mathrm{old}}(x)}$ denotes the probability ratio between the policy under optimization and the policy for sampling, $A(x)$ is the advantage function, and $\epsilon$ is the clipping parameter. The second term, $\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon)$, modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $\gamma(\theta)$ outside of the interval $[1-\epsilon,1+\epsilon]$. Drawing samples from $p_{\mathrm{old}}(x)$ rather than $p_{\theta}(x)$ stems from the importance sampling and reduce the number of costly evaluation of the reward function.
@@
Appendix D Dielectric constant prediction model
The total dielectric constant is defined as
$\varepsilon_{\mathrm{total}}=\varepsilon_{\mathrm{ionic}}+\varepsilon_{\mathrm{elec}},$ (S4)
where $\varepsilon_{\mathrm{ionic}}$ and $\varepsilon_{\mathrm{elec}}$ denote the ionic and electronic contributions to the dielectric constant, respectively. In fact, the dielectric constant is fundamentally a tensor $\varepsilon_{ij}$, and the scalar dielectric constant $\varepsilon_{\mathrm{total}}$ is the mean of the eigenvalues of the total dielectric tensor. The dielectric tensor can be written as
$\varepsilon=\begin{bmatrix}\varepsilon_{11}&\varepsilon_{12}&\varepsilon_{13}\\ \varepsilon_{21}&\varepsilon_{22}&\varepsilon_{23}\\ \varepsilon_{31}&\varepsilon_{32}&\varepsilon_{33}\\ \end{bmatrix},$ (S5)
where $\varepsilon_{12}=\varepsilon_{21}$, $\varepsilon_{13}=\varepsilon_{31}$, and $\varepsilon_{23}=\varepsilon_{32}$. While the dielectric tensor is characterized by 9 components, only up to 6 are independent elements. Moreover, this number decreases with higher material symmetry [80].
We use the M3GNet [30] model implemented in the MatGL [81] repository to predict the 6 independent components of the total dielectric tensor using the DFPT calculations in Materials Project dataset [45, 46]. The 6 independent components were predicted by the model and assembled into a $3\times 3$ matrix, from which the total dielectric constant was determined by solving for its eigenvalues. Following the data processing in Ref. [82], we clean up the dataset by restricting the components in total dielectric constants $\varepsilon_{ij}\in[-10,100]$ for $\forall i,j$, which resulted in a dataset of $\sim$ 7,000 structures. The above dataset is then divided into training, validation, and test subsets with a ratio of 8:1:1, respectively, for the purpose of experimental analysis. The dielectric property prediction model is trained with the $L1$ loss and batch size of 64 and learning rate of 1e-4, using the Adam optimizer with default parameters. Other training settings remained as the default configuration of MatGL.
Figure S4: Performance of M3GNet on the test set of our refined MP dataset for $\varepsilon_{\mathrm{total}}$.
Figure S4 illustrated the performance of M3GNet on the by calculating the total dielectric constant. It can be observed that despite constraining the data range, the model’s generalization within this range remains limited, often underestimating the total dielectric constant of the materials. This issue can be resolved by implementing a more robust prediction model or by expanding the dataset to include a broader range of dielectric constants.
Appendix E Dielectric materials discovered by CrystalFormer-RL
As summarized in the Table LABEL:table:dielectric_samples, the reinforcement fine-tuned model generates a series of stable dielectric materials with high FoM values, with only two samples present in the DFPT dataset. The labels of the LiTl2InF6 compound are band gap $4.10$ eV and total dielectric constant $27.49$. The labels of Ba2TaFeO6 are band gap $1.86$ eV and total dielectric constant $46.92$. Both are close to the predicted values in Table LABEL:table:dielectric_samples.
Table S2: Dielectric materials discovered by CrystalFormer-RL, ranked according to their figures of merit (FoM): $\varepsilon_{\mathrm{total}}\cdot E_{g}$. See details of the samples at https://drive.google.com/file/d/1ouz0RRgntiSoVRe5VP_cemnAdcb3MpTj/view?usp=sharing.
Formula $E_{g}$ (eV) 1 $\varepsilon_{total}$  1 FoM (eV) $E_{\mathrm{hull}}$ (eV/atom) 2 Not in Alex-20 Not in DFPT dataset
LiNdTl2F6 4.54 41.73 189.54 0.09
LiLaTl2F6 4.48 40.44 181.35 0.08
LiErTl2F6 4.73 36.56 172.78 0.05
LiScTl2F6 4.40 37.41 164.57 0.03
KDyTl2F6 6.32 23.49 148.50 0.05
NaTbTl2F6 5.02 29.17 146.56 0.02
LiYTl2F6 4.69 30.98 145.27 0.04
LiTmTl2F6 4.53 29.19 132.27 0.03
LiNiPb2F6 4.12 30.43 125.26 -0.04
Li6PbCl8 4.64 25.19 116.76 0.08
Cs2BaSrF6 5.38 21.32 114.75 0.04
Cs2SmTlF6 5.69 20.12 114.50 0.02
LaMnTl2F6 3.03 36.01 109.15 -0.04
Cs2SrCaF6 6.10 17.86 108.92 0.01
Cs2LiNdF6 6.44 16.90 108.90 0.06
ScTl2WO6 1.99 54.47 108.50 0.10
ScPb2WO6 1.53 70.52 107.90 0.10
Ba2TaMnO6 2.75 39.00 107.37 0.08
CsBaF3 5.81 18.38 106.78 0.08
LiTl2InF6 3.72 28.74 106.76 0.03
Cs2LiPrF6 6.52 16.25 105.99 0.07
Rb2NdTlF6 5.29 20.02 105.90 0.09
Ba2TaFeO6 2.07 50.19 103.81 0.02
Cs2ScTlF6 4.87 21.26 103.51 0.00
Cs2LiLaF6 5.96 17.35 103.50 0.07
Cs2KNdF6 5.96 17.37 103.49 0.01
Cs2RbNdF6 6.40 16.16 103.49 0.02
Cs2BaCaF6 6.07 17.00 103.13 0.04
Cs2LuTlF6 5.48 18.61 101.96 0.00
Cs2BaPbF6 4.66 21.82 101.69 0.04
Cs2RbSrF6 5.89 17.22 101.47 0.09
Cs3NdF6 6.18 16.33 100.95 0.04
  • 1
    The band gap and total dielectric constant are predicted by machine learning models.
  • 2
    The energy above the convex hull is computed via the Orb model based on Alexandria convex hull.
Table S3: Dielectric materials discovered in the Alex-20 dataset. Superscript ’eq’ denotes the prediction results for equilibrium structure in the Alex-20 dataset
Formula $E_{g}$ (eV) 1 $\varepsilon_{total}$  1 FoM (eV) $a$ ($\AA$2 $E_{g}^{eq}$ (eV) 1 $\varepsilon_{total}^{eq}$  1 FoMeq (eV) $a^{eq}$ ($\AA$)  2
LiNdTl2F6 4.54 41.73 189.54 6.23 4.69 41.92 196.41 6.23
LiLaTl2F6 4.48 40.44 181.35 6.31 4.63 37.72 174.53 6.22
LiErTl2F6 4.73 36.56 172.78 6.09 4.71 36.93 173.98 6.11
LiScTl2F6 4.40 37.41 164.57 6.00 4.54 36.24 164.49 6.00
LiYTl2F6 4.69 30.98 145.27 6.16 4.70 30.31 142.52 6.13
LiTmTl2F6 4.53 29.19 132.27 6.09 4.59 29.99 137.59 6.09
Cs2SmTlF6 5.69 20.12 114.50 7.11 5.73 19.84 113.69 7.01
Cs2SrCaF6 6.10 17.86 108.92 6.73 6.04 16.59 100.17 6.64
Cs2LiNdF6 6.44 16.90 108.90 6.41 6.41 16.82 107.89 6.41
Ba2TaMnO6 2.75 39.00 107.37 5.85 1.88 35.08 65.82 5.66
CsBaF3 5.81 18.38 106.78 7.34 5.81 18.94 109.98 5.14
LiTl2InF6 3.72 28.74 106.76 6.06 3.90 26.52 103.32 6.05
Cs2LiPrF6 6.52 16.25 105.99 6.46 6.43 15.79 101.48 6.43
Cs2ScTlF6 4.87 21.26 103.51 6.82 4.86 20.78 100.99 6.73
Cs2LiLaF6 5.96 17.35 103.50 6.45 5.68 16.72 94.88 6.44
Cs2KNdF6 5.96 17.37 103.49 7.03 5.81 14.85 86.31 6.90
Cs2RbNdF6 6.40 16.16 103.49 7.11 6.45 15.85 102.16 7.04
Cs2BaCaF6 6.07 17.00 103.13 6.87 6.08 16.77 101.99 6.84
Cs2BaPbF6 4.66 21.82 101.69 7.14 4.64 21.61 100.23 7.10
Cs2RbSrF6 5.89 17.22 101.47 7.05 5.41 17.39 94.01 7.07
Cs3NdF6 6.18 16.33 100.95 7.29 5.18 13.51 69.98 7.22
  • 1
    The band gap and total dielectric constant are predicted by machine learning models.
  • 2
    The lattice constant is measured using the primitive cell.

### v2

@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Many of the generated samples are present in the Alex-20 dataset which is in line with the S.U.N ratio 5.1% of the base model on space group No.225. This nevertheless demonstrates that the reinforcement fine-tuning procedure can guide the generative process toward discovering candidates within the pretrained dataset with no property label. Notably, two materials in Table LABEL:table:dielectric_samples are absent from the Alex-20 dataset, highlighting the ability of RL to guide the generative process toward novel samples beyond the pretrained dataset.
Table 1: Dielectric materials discovered by CrystalFormer-RL, ranked according to their figures of merit (FoM): $\varepsilon_{\mathrm{elec}}\cdot E_{g}$. The band gap, dielectric constant, and energy above the convex hull are all calculated by DFT. The two highlighted rows are not in the Alex-20 dataset. See details of the samples at https://drive.google.com/file/d/1bEGzeDPybiDZw5lCzLwRapoMSbaGfQN0/view?usp=sharing.
Formula $E_{g}$ (eV) 1 $\varepsilon_{elec}$ FoM (eV) $E_{\mathrm{hull}}$ (eV/atom)
Cs2LiLuF6 7.37 2.38 17.52 0.09
Cs2LiTmF6 7.35 2.37 17.40 0.08
Cs2LiHoF6 7.32 2.37 17.39 0.09
Cs2LiErF6 7.29 2.37 17.28 0.09
Cs2LiTbF6 7.21 2.36 17.05 0.09
Cs2LiDyF6 7.17 2.34 16.80 0.09
CsCaF3 6.94 2.32 16.10 0.06
Cs2LiScF6 6.64 2.40 15.92 0.09
Cs2NaTmF6 6.98 2.27 15.84 0.05
Cs2NaErF6 6.96 2.27 15.77 0.05
Cs2NaLuF6 6.93 2.26 15.63 0.05
Cs2NaHoF6 6.90 2.26 15.59 0.05
Cs2NaYF6 6.89 2.25 15.48 0.03
Cs2LiLuCl6 5.34 2.90 15.48 0.01
Rb2LiLuF6 7.01 2.19 15.34 0.07
Cs2LiHoCl6 5.16 2.86 14.78 0.01
Cs2LiTbCl6 5.07 2.89 14.66 0.01
Y2ErF6 7.15 1.96 14.01 0.05
Cs2SrMgCl6 5.07 2.74 13.89 0.03
Cs2MgPbF6 5.29 2.60 13.77 0.08
K2Rb2AlF6 6.40 2.02 12.92 0.06
Cs2RbDyCl6 4.92 2.46 12.09 0.03
Cs2ErGaCl6 3.82 3.05 11.64 0.04
Cs2BaPbI6 3.17 3.62 11.48 0.07
Cs2NaAlCl6 3.97 2.77 11.00 0.03
Rb2TmGaCl6 3.60 2.99 10.77 0.06
We compare the DFT-calculated band gap and electronic dielectric constant of the generated samples with the predictions from the MEGNet and AnisoNet models. As shown in Figure 7, the ML predictions align well with the DFT results, validating the effectiveness of the RL fine-tuning approach for property-guided material generation.
Figure 7: ML predicted vs. DFT calculated (a) energy above the convex hull $E_{\mathrm{hull}}$, (b) band gap $E_{g}$ and (c) electronic dielectric constant $\varepsilon_{\mathrm{elec}}$ for the discovered dielectric materials. Panels (b) and (c) include only materials with $E_{\mathrm{hull}}<0.1$ eV/atom according to the DFT calculation. The two stars indicate the same two novel samples as Fig. 6(a).
@@ IV Discussions
Reinforcement fine-tuning of materials generative model is useful for optimizing other materials properties such as the thermoelectric figure of merit, thermal conductivity, and superconducting critical temperature. In addition to a foundational generative model such as the CrystalFormer for the materials prior $p(x)$, the framework requires a reliable ML prediction model for the likelihood $p(y|x)$, which is a research topic with rapid progress both on generalizability and accuracy, see, e.g. [54, 55, 56].
@@
It is often easier to discriminate the utility of a given material than to generate a material with the desired utility. Reinforcement fine-tuning of a materials generative model using discriminative feedback offers a way to bridge such discrimination-generation gap. RL on the autoregressive transformer such as CrystalFormer [1] is particularly convenient given a variety of post-training techniques available for LLM. Although those RL techniques are also applicable to natural language based materials generative models [57, 58, 5], we believe finetuned CrystalFormer is more advantageous because it natively speaks the language of crystalline materials, which cherishes the fundamental inductive bias of crystalline materials such as space group symmetry and periodicity. Furthermore, the idea of reinforcement finetuning can also be applied to the material generative models based on diffusion [59, 60]. The diffusion model is typically fine-tuned through the incorporation of auxiliary networks to condition the generation on external inputs [61], while RL eliminates the necessity for these additional network components. For example, Refs. [62, 63] employ the RL algorithm to fine-tune the pre-trained text-to-image diffusion model, leading to improved sample quality and diversity. In the mean time, applying RL with KL regularization to the diffusion model is challenging due to the need for precise log-likelihood calculations [64].
@@
It is also interesting to compare active learning and RL in the context of generative materials discovery. Active learning [65, 66] selects crystal samples back to the generative training dataset, which is similar to the SFT approach discussed in Sec. III.1. Similar to pretrain, the active learning procedure optimizes the forward KL divergence between the empirical distribution of the training dataset and the generative model. In contrast, RL minimizes the reverse KL divergence [25] between the generative model and the target distribution Eq. (2). There are both empirical and theoretical evidences that the RL approach delivers better performance compared to the active learning approach in the contexts of LLM fine-tuning [67, 68]. Consequently, the primary advantage of RL over active learning is its ability to directly incorporate task-specific reward signals, enabling the model to generate outputs that more effectively meet practical application requirements. Moreover, we suspect the difference in the mode covering versus mode seeking in the active learning versus RL may also be a reason for the difference in the performance [69, 26].
@@
Figure 7 compares the ML-predicted and DFT-calculated results, revealing that while the overall trends are consistent, noticeable discrepancies remain between the predictions and the DFT evaluations. This reveals a fundamental challenge in the application of RL in materials science: what is the verifiable ground truth reward signal. Given the intricacies involved in material synthesis and characterization [71], such a challenge is particularly concerning as faulty reward functions [72] which fails to align with the final goal will suggest structures with high reward but fail for the final experiment or even intermediate DFT verifications. We believe the design of reward functions that are both actionable and well-aligned with experimental outcomes, with the insights of human experts, is crucial for ensuring the reliability of the present and many related efforts.
@@
This study focuses on how generative models can be enhanced and specialized through the integration of property prediction models. However, the reverse question: how property prediction models can be improved by leveraging generative models—is also an intriguing question worth future investigation. For example, the Orb models employ a two-stage training framework that integrates generative and supervised learning to achieve state-of-the-art performance [13]. Initially, the model is pre-trained as denoising diffusion models on datasets of stable materials. The pre-trained model then serves as the initialization for an MLIP, which is subsequently fine-tuned in a supervised manner to predict the energy, forces, and stress of materials. This approach exemplifies how generative pre-training can enhance the performance of discriminative models such as MLIP. Extending this idea further, Ref. [12] has demonstrated that fine-tuning pretrained MLIP also gives better performance for property prediction models.
@@ Acknowledgements.
We thank Han Wang, Linfeng Zhang, Jian Lv, Hongteng Xu, Shigang Ou, Xiaoshan Luo, Mingzhang Yang, Tianping Ying, Ling Lu, Weiluo Ren and Huaxian Jia for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, and No. T2121001, Cross-Disciplinary Key Project of Beijing Natural Science Foundation No. Z250005, National Key Projects for Research and Development of China Grant. No. 2021YFA1400400, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000.
@@ Appendix A The CrystalFormer model card
The base model was trained for 5500 epochs with a total batch size of 8000 over eight A100 GPUs using the Adam optimizer [77]. The learning rate was set to 8e38e-3.
The reinforcement fine-tuning was performed for 100 - 200 steps, depending on the task. For the reinforcement fine-tuning, we use a total batch size of 1000 over two A100 GPUs and the Adam optimizer with a learning rate of 1e41e-4. The pre-training phase required approximately 1000 GPU hours, while the reinforcement fine-tuning phase took around 12 to 24 GPU hours.
For the supervised fine-tuning (SFT) of the CrystalFormer, we generate 10,000 crystal samples and relax them using the Orb models. Only those stable samples with $E_{\mathrm{hull}}<0.1$ eV/atom were retained for SFT. The SFT was performed for 55 epochs when validation loss stopped improving, with a batch size of 100 using the Adam optimizer with a learning rate of 1e51e-5 on a single A100 GPU.
@@
   Hidden layer dimension for the Wyckoff letter of the first atom    256
@@
   RL regularization coefficient $\tau$    0.1    2 for FoM reward
@@ Appendix B Training dataset
Figure S3: Composition and structure validity of samples generated by the pretrained CrystalFormer using the MP-20 and Alex-20 dataset, as well as reinforcement fine-tuned model using the energy above hull as the reward function.
@@ Appendix C Details of reinforcement fine-tuning implementation
where $\gamma(\theta)=\frac{p_{\theta}(x)}{p_{\mathrm{old}}(x)}$ denotes the probability ratio between the policy under optimization and the policy for sampling, $A(x)$ is the advantage function, and $\epsilon$ is the clipping parameter. The second term, $\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon)$, modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $\gamma(\theta)$ outside of the interval $1-\epsilon,1+\epsilon]$. Drawing samples from $p_{\mathrm{old}}(x)$ rather than $p_{\theta}(x)$ stems from the importance sampling and reduce the number of costly evaluation of the reward function.
@@
Appendix D Computational details
D.1 Machine learning force field
@@
D.2 DFT calculations
The DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [46] and all-electron projector-augmented wave method [84], as implemented in the VASP code [85]. All parameters of the calculations including settings of PBE functional, Hubbard U corrections, and ferromagnetic initialization are chosen to be consistent with Materials Project by using of MPRelaxSet function in pymatgen [83]. All structures containing Yb element are ignored when calculating energy above the hull due to they are unavailable from the Materials Project at the time of writing. The DFPT input files are generated using the MPStaticSet in pymatgen, with a k-point density of 3,000 per reciprocal atom and a plane-wave energy cutoff of 600 eV, consistent with the Materials Project settings [45].